# Victoria, BC, Canada

### v2

@@ Abstract
Mixed methods research is often used in software engineering, but researchers outside of the social or human sciences often lack experience when using these designs. This paper provides guidelines and advice on how to design mixed method research, and to encourage the intentional, rigorous, and innovative use of mixed methods in software engineering. It also presents key characteristics of core mixed method research designs. Through a number of fictitious but recognizable software engineering research scenarios, we showcase how to choose suitable designs and consider the inevitable trade-offs any design choice leads to. We describe several antipatterns that illustrate what to avoid in mixed method research, and when mixed method research should be considered over other approaches.
@@
Mixed methods Research methods Methodology Guidelines
@@ 1 Introduction
Software engineering (SE) researchers use a variety of research methods, often mixing and matching them. Despite the widespread prevalence of mixing methodsbryman_integrating_2006 ; creswell_best_2011 ; poth_innovation_2018 ; ladner_mixed_2019 ; poth_sage_2023 , there is little guidance in SE on how to do so systematically, in a way that is well motivated, carefully designed, robustly executed, and leads to additional benefits not otherwise achievable from the use of a single research method alone. Mixing methods is particularly pertinent in today’s complex socio-technical SE contexts storey_who_2020 ; hoda2024Book where quantitative and qualitative methods, when combined, can bring insights on the nuanced nature of the problem or solution being studied.
Although mixed methods are known in SE, there is not enough debate on the methodological issues, challenges and trends that set the current scenario of mixed methods research in SE. Di Penta and Tamburri Dipenta discuss how mixed methods in empirical SE combine quantitative and qualitative research, and how qualitative methods can enhance quantitative studies and vice versa. Critically, there is no clear guidelines explaining mixed methods research for SE. To the best of our knowledge, our paper is the first to provide guidelines on mixed methods research designs in an SE context that researchers can adopt when designing studies that foresee mixing qualitative (quaL) and quantitative (quaN) methods.
Our paper is aimed at researchers unfamiliar with mixed methods (the use of both quaN and quaL methods) and those keen to improve their understanding and practice. We share key concepts and provide guidance on how to mix and match methods, presenting different characteristics of mixed methods and some common designs.
@@
The timing of this paper is important, as SE researchers today face a complex world where software is used to address complex problems with wide-scale societal impact. There is also a strong drive for innovation in the face of rapidly changing technologies and practices, all occurring in socio-technical contexts storey_who_2020 ; hoda_socio-technical_2022 . The increasing interest in studying the adoption of generative AI technologies in everyday SE practice is but one latest example of the constantly changing SE research landscape bano2023LLM . Choosing which mixed methods to use and knowing how to use them effectively is challenging, and is often overlooked in SE research and education curricula and thus, experience is lacking in how to use these methodological designs.
@@
Inspired by a previous paper, one of us co-authored on selecting empirical methods for SE research easterbrook_selecting_2008 , this paper includes practical illustrative examples in the form of four fictional SE researchers and research scenarios, representing common mixed method contexts, designs, and outcomes. These four research scenarios are used to exemplify and motivate the use of mixed method research designs throughout the paper. While not exhaustive, these scenarios not only help to guide design choice but also emphasize the trade-offs faced in making some common mixed method study design decisions. As well as showcasing design choices for a variety of scenarios, the paper also calls out best practices and a number of antipatterns that should be avoided.
@@
The rest of this paper is organized as follows. In Section 2, we present a brief background of the most prominent and relevant influences in mixed methods research design over the decades. In Section 3, we present four core principles that underlie mixed method research: methodological rationale, novel integrated insights, procedural rigor and ethical research. In Section 4, we discuss some notable characteristics of mixed method designs, as well as common mixed methods research designs. In Section 5, we present four scenarios of research studies that fictional researchers pursue. We describe the methodological choices these characters make and how they followed the four core principles mentioned in Section 3. In Section 6, we use the concept of antipatterns to present what not to do in mixed method research. In Section 7, we delve in to the importance of integration in mixed methods research, discuss mixed methods research designs with other research designs (such as multi-method, case studies and secondary studies), and share a call to action before concluding the paper in Section 8. This paper includes an Appendix which contains a glossary of terms we use throughout the paper for ease of reference.
@@ 2 A Brief Overview of Mixed Methods Research
It is only in the mid-20th century, with roots in disciplines such as sociology and anthropology, that mixed methods research (MMR) gained traction with the formalization of approaches by authors including Creswell and Plano Clark creswell_designing_2018 . According to these authors, “The formative period of mixed methods began in the 1950s and continued up until the 1980s. This period saw the initial interest in using more than one method in a study” but these authors did not explicitly identify these approaches as mixed methods. Furthermore, the different typologies of mixed method designs did not emerge until the 1990s maxwell_expanding_2016 ; Greene1989 .
@@
In the social sciences field, Campbell and Fisk’s paper in 1959 campbell_convergent_1959 is recognized as the first adoption of multiple research methods for validation purposes. They also defined this design as multi-operationalism since the authors propose a multitrait, multimethod matrix that intersects quaL-quaN methods with a specific set of traits as a means for enforcing validity of research conclusions. In the mid 1960s, Webb et al. webb_unobtrusive_1966 introduced the concept of triangulation, followed by Denzin in the 1970s denzin_research_1978 who detailed how to triangulate using both within-methods triangulation and between-methods triangulation, emphasizing the latter as a more powerful means to overcome bias and as a way to pragmatically mix methods.
@@
Although the combination of quaL-quaN methods has been used in practice for many years, the first evidence of a more formal use of mixed methods is identifiable only in the late 1980s, discussed by Guest and Fleming guest_mixed_2015 and Johnson et al. johnson_toward_2007 who coined mixed methods research as the third paradigm, coexisting with quaL and quaN studies. The common denominator in these early conceptualizations of mixed methods is that MMR combines quaL-quaN elements. It is only recently that researchers have turned their attention to proposing typologies of mixed methods, i.e., which describe how to structure and design a MMR by combining quaL and quaN elements according to some criteria as stated by Teddlie and Tashakkori teddlie_foundations_2009 .
@@
This rise of formalized mixed method approaches has led to numerous exciting discussions, but often with different design representations and overlapping terms. Indeed, not even the basic definition of what can be qualified as an MMR study sees consensus (see Section 1 in  poth_sage_2023 ). Part of this issue originates in the “paradigm war”, where some researchers have polarized preferences for using quaN (e.g., positivism) or quaL (e.g., constructivism) data methods. It is not our goal to review these research paradigms, design typologies or resolve paradigm debates in this paper, but it is essential to note that some research paradigms that underpin MMR, such as pragmatism, emphasize the importance of using the most effective methods to answer research questions. Moreover, the philosophical foundation of a researcher’s choice will inevitably impact how quaL and quaN approaches are integrated. Today, MMR is widely recognized and utilized across diverse fields such as psychology, education, public health, and management, and these approaches offer a flexible and robust framework for addressing multifaceted research inquiries and advancing knowledge in complex socio-technical domains poth_innovation_2018 ; ladner_mixed_2019 .
@@
With the increasing complexity of MMR designs, the problem of discussing and visually presenting complex designs to readers became apparent. Consequently, as noted by Creswell and Plano Clark (poth_sage_2023, , Chapter 2) when they revisited MMR designs twenty years after their important chapter was first published in 2003 tashakkori_handbook_2003 , the most recent studies decreased their focus on the quaN and quaL components and increased attention towards the integration processes across design types. To reflect this change about the central elements of the MMR designs, Figure 1 represents the simplification of the interconnection of the core components of a MMR design. From this perspective, treating the data beyond quaL and quaN analysis and gaining insight through integration and meta-inferences become central.
@@
To conclude our brief overview of MMR, current trends in MMR can be listed as a call for openness and creativity during the research process, integrating mixed methods in interdisciplinary research, and the use of digital tools for data collection and analysis. Furthermore, researchers will be increasingly challenged to keep pace with a rapid evolution in design terminology and naming practices. Given such dynamism, paying attention to the many sources of contextual influences on the design of MMR becomes paramount poth_sage_2023 . Recognizing the challenges these trends impose on researchers, we provide practical design guidance of MMR for SE researchers in our paper. To start, in the next section, we introduce the definition of MMR we use and then we describe four fundamental principles to be followed when using MMR.
@@ 3 Principles to Guide Mixed Method Research
Building on Poth poth_innovation_2018 , we define mixed methods research (MMR) as, a research approach where multiple procedures are used to collect, analyze, and integrate data to address a research problem. For mixed methods, both qualitative and quantitative data are integrated to produce novel insights.
Poth poth_innovation_2018 also presents four core guiding principles for MMR that we have adapted slightly to be applicable to SE (see Figure 2). Two of the principles, methodological rationale and novel integrated insights, relate to understanding why we should consider using a MMR design. The other two, procedural rigor and ethical research, relate to understanding how we should go about designing and applying MMR. While some aspects of these principles apply to single method applications and can be considered good research practice in general, others are particularly relevant to MMR and generally more complex in that context. We delve into these four principles below.
Figure 2: Principles for Mixed Method Research Designs, as inspired by Poth poth_innovation_2018 .
@@ 3.1 Methodological Rationale
Principle of Methodological Rationale: Mixed method research study designs should explain how the mixed research design and data procedures are relevant for addressing a given research problem. By adhering to this principle, an MMR study description can address the question why did we select a particular MMR research design and set of procedures?
There are several conceptual reasons for using mixed methods [^1] The reasons we list are inspired by or earlier proposed by Ladner ladner_mixed_2019 , Bryman bryman_social_2012 , Creswell creswell_designing_2018 and McGrath mcgrath_methodology_1995 . The rationale for the methodology may be one or more of the following:
  • Complementarity: one procedure helps to deepen, confirm or enhance the insights of another procedure ladner_mixed_2019 ; bryman_social_2012 .
  • Expansion: one procedure is used to ask different questions that are apparent up front or emerge throughout the study (in parallel or sequentially) ladner_mixed_2019 ; bryman_social_2012 .
  • Development: one procedure is used to inform or improve another procedure (e.g., interviews may be used to inform the development of a questionnaire) ladner_mixed_2019 ; bryman_social_2012 .
  • Triangulation: a different type of procedure is used to corroborate earlier insights leveraging multiple quality criteria afforded by use of different procedures mcgrath_methodology_1995 notably generalizability, realism, precision, and control.
  • Credibility: another procedure may be used to increase the integrity and truth of the findings from a former procedure bryman_social_2012 . Mixed methods can increase credibility by bringing multi-dimensional philosophical and theoretical lenses to the findings mertens_mixed_2013 .
  • Explanation: one procedure can be used to resolve previous contradictory, surprising or inconclusive findings bryman_integrating_2006 .
@@ 3.2 Novel Integrated Insights
Principle of Novel Integrated Insights: Mixed methods research studies should generate new insights through the integration of the procedures, which in turn addresses the reason(s) for the use of mixed methods. An MMR study report should clarify the novelty of the study findings and explain how methods and/or data were integrated. By adhering to this principle, an MMR study report can address the question what did we gain from using MMR?
Succinctly, an MMR study should deliver on its promise of additional benefits to what could have been afforded by using a single method:
@@
  • Explaining unexpected results: where the use of one procedure produced unexpected results, another can be employed in a bid to further explore why those unexpected results were encountered.
@@ 3.3 Procedural Rigor
Principle of Procedural Rigor: Mixed method studies should apply the study procedures rigorously and explain them effectively. By adhering to this principle, an MMR study report can address the question how well did we conduct the study?
Rigor in any research study is always a critical concern, and it is often a difficult concept for students to know how to assess. Rigor is particularly challenging for mixed methods. MMR studies should first adhere to the rigor of the individual studies, but there are also some concerns that apply to the “mixing” aspect. For this principle, we are guided by a mixed methods rigor appraisal tool called MMAT (Mixed Methods Appraisal Tool) oliveira_mixed_2021 . The following are posed as questions to be considered when self-reflecting on the rigor of a mixed design:
@@
  • Are the different methods used rigorously conducted? MMR studies should adhere to the rigor of the individual studies used in the mixed study (referring to the appropriate quality criteria for each study [^4] See the ACM Empirical Standards sigsoft ).

### v3

@@ Abstract
Mixed methods research is often used in software engineering, but researchers outside of the social or human sciences often lack experience when using these designs. This paper provides guiding principles and advice on how to design mixed method research, and to encourage the intentional, rigorous, and innovative application of mixed methods in software engineering. It also presents key properties of core mixed method research designs. Through a number of fictitious but recognizable software engineering research scenarios, we showcase how to choose suitable mixed method designs and consider the inevitable trade-offs any design choice leads to. We describe several antipatterns that illustrate what to avoid in mixed method research, and when mixed method research should be considered over other approaches.
@@
Mixed methods Research methods Methodology Guiding Principles
@@ 1 Introduction
Software engineering (SE) researchers use a variety of research methods, often mixing and matching them. Mixing methods is particularly pertinent in today’s complex socio-technical SE contexts storey_who_2020 ; hoda2024Book where quantitative and qualitative methods, when combined, can bring insights on the nuanced nature of the problem and/or solution being studied. Before discussing the use and state of mixed methods in SE research, we first share the definition of mixed methods research we use throughout our paper [^1] This definition is in line with the definitions by other scholars of mixed methods research, such as  poth_sage_2023 .:
Mixed methods research (MMR) is a research approach where multiple methods are used to collect, analyze, and integrate both qualitative and quantitative data to address a research problem and produce novel insights.
There is widespread prevalence of mixing methods in the social sciences bryman_integrating_2006 ; creswell_best_2011 ; poth_innovation_2018 ; ladner_mixed_2019 ; poth_sage_2023 and well developed guidelines for mixed methods research in the related discipline of Information Systems (IS)  venkatesh2013bridging ; venkatesh2016guidelines . However, while mixed methods are recognized and utilized in Software Engineering (SE), there is limited discussion on the unique methodological issues, challenges, and emerging trends shaping mixed methods research in this domain. Di Penta and Tamburri Dipenta explore the integration of quantitative and qualitative approaches in empirical SE, highlighting how each can complement the other. However, there remains a gap in providing actionable guidance for applying mixed methods research specifically within SE. This paper aims to fill that gap by offering practical guiding principles to support researchers in designing studies that effectively and rigorously combine qualitative (quaL) and quantitative (quaN) methods in an SE context.
@@
Our paper is aimed at researchers unfamiliar with mixed methods and those keen to improve their understanding and practice. We share key concepts and provide guidance on how to mix and match methods, presenting different properties of mixed methods and some common designs. The timing of this paper is important, as SE researchers today face a complex world where software is used to address complex problems with wide-scale societal impact. There is also a strong drive for innovation in the face of rapidly changing technologies and practices, all occurring in socio-technical contexts storey_who_2020 ; hoda_socio-technical_2022 . The increasing interest in studying the adoption of generative AI technologies in everyday SE practice is but one latest example of the constantly changing SE research landscape bano2023LLM . Choosing which mixed methods to use and knowing how to use them effectively is challenging, and is often overlooked in SE research and education curricula and thus, experience is lacking in how to use these methodological designs.
@@
Inspired by a previous paper one of us co-authored on selecting empirical methods for SE research easterbrook_selecting_2008 , this paper includes practical illustrative examples in the form of four fictional SE researchers and research scenarios, representing common mixed method contexts, design decisions, and outcomes. These four research scenarios are used to exemplify and motivate the use of mixed method research designs throughout the paper. While not exhaustive, these scenarios not only help to guide design choice but also emphasize the trade-offs faced in making some common mixed method study design decisions. As well as showcasing design choices for a variety of scenarios, the paper also calls out best practices and a number of antipatterns that should be avoided.
Through this paper, we aim to demystify, clarify, and define MMR and its various properties, designs, and application scenarios for the SE research community. We encourage and call for SE researchers to use mixed methods and to do so in a more principled way. Our hope is that it will serve as the start of a robust conversation in the community about intentional, rigorous, and innovative use of MMR in SE, one that inevitably others will build upon in the future.
@@
The rest of this paper is organized as follows. In Section 2, we present a brief background of the most prominent and relevant influences in mixed methods research design over the decades. In Section 3, we present four core principles that underlie mixed method research: methodological rationale, novel integrated insights, procedural rigor and ethical research. In Section 4, we discuss some notable properties of mixed method designs, as well as common mixed methods research designs. In Section 5, we present four scenarios of research studies that fictional researchers pursue. We describe the methodological choices these characters make and how they followed the four core principles mentioned in Section 3. In Section 6, we use the concept of antipatterns to present what not to do in mixed method research. In Section 7, we discuss mixed methods research designs with other research designs (such as multi-method, case studies and secondary studies) and share a call to action before concluding the paper in Section 8. This paper includes an Appendix which contains a glossary of terms we use throughout the paper for ease of reference.
@@ 2 A Brief Overview of Mixed Methods Research
It is only in the mid-20th century, with roots in disciplines such as sociology and anthropology, that mixed methods research (MMR) gained traction with the formalization of approaches by authors including Creswell and Plano Clark creswell_designing_2018 . According to these authors, “The formative period of mixed methods began in the 1950s and continued up until the 1980s. This period saw the initial interest in using more than one method in a study” but these authors did not explicitly identify these approaches as mixed methods. Furthermore, the different typologies of mixed method designs did not emerge until the 1990s maxwell_expanding_2016 ; Greene1989 . In the social sciences field, Campbell and Fiske’s paper from 1959 campbell_convergent_1959 is recognized as the first adoption of multiple (that included mixed) research methods for validation purposes. They defined the multiple method design as multi-operationalism since the authors propose a multitrait, multimethod matrix that intersects quaL-quaN methods with a specific set of traits as a means for enforcing validity of research conclusions. In the mid 1960s, Webb et al. webb_unobtrusive_1966 introduced the concept of triangulation, followed by Denzin in the 1970s denzin_research_1978 who detailed how to triangulate using both within-methods triangulation and between-methods triangulation, emphasizing the latter as a more powerful means to overcome bias and as a way to pragmatically mix methods.
@@
Although the combination of quaL-quaN methods has been used in practice for many years, the first evidence of a more formal use of mixed methods is identifiable only in the late 1980s Kuhn1977 , as discussed by Guest and Fleming guest_mixed_2015 and Johnson et al. Johnson2004 ,johnson_toward_2007 who point out how mixed methods research combines qualitative and quantitative approaches, forming a third paradigm that complements traditional methods to yield more comprehensive, balanced, and insightful results. The common denominator in these early conceptualizations of mixed methods is that MMR combines quaL-quaN elements. It is only recently that researchers have turned their attention to proposing typologies of mixed methods, i.e., which describe how to structure and design a MMR by combining quaL and quaN elements according to some criteria as stated by Teddlie and Tashakkori teddlie_foundations_2009 .
@@
This rise of formalized mixed method approaches has led to numerous discussions, but often with different design representations and overlapping terms. Indeed, not even the basic definition of what can be qualified as an MMR study sees consensus (see Section 1 in  poth_sage_2023 ). Part of this issue originates in the “paradigm war”Bryman2008 where Bryman discusses the historical association of quaN research with positivism and quaL research with constructivism. He argues that these conflicts have diminished over time, giving way to a more pragmatic approach that embraces methodological pluralism as researchers are increasingly combining quantitative and qualitative methods to enrich their studies, moving beyond rigid adherence to a single paradigm. This shift reflects a growing acceptance of mixed methods as a valuable research strategy. It is not our goal to review these research paradigms, design typologies or resolve paradigm debates in this paper, but it is essential to note that some research paradigms that underpin MMR, such as pragmatism, emphasize the importance of using the most effective methods to answer research questions. Moreover, the philosophical foundation of a researcher’s choice will inevitably impact how quaL and quaN approaches are integrated.
Today, MMR is widely recognized and utilized across diverse fields such as psychology, education, public health, information systems and management, and these approaches offer a flexible and robust framework for addressing multifaceted research inquiries and advancing knowledge in complex socio-technical domains poth_innovation_2018 ; ladner_mixed_2019 . In software engineering, we can learn much from related disciplines, such as information systems venkatesh2013bridging ; venkatesh2016guidelines , that articulate principles and guidelines of how to conduct mixed methods research. The presented guidelines by Venkatesh et al. venkatesh2013bridging , building on top of the mixed methods research by Tashakkori and Creswel TashakkoriCreswell2008 , Teddlie and Tashakkori teddlie2003major ; teddlie_foundations_2009 , Creswell creswell_research_2003 among others, clearly articulate that “the decision to conduct mixed methods research should hinge on the research question, purpose, and context.”
@@
With the increasing complexity of MMR designs, the problem of discussing and visually presenting complex designs to readers became apparent. Consequently, as noted by Creswell and Plano Clark (poth_sage_2023, , Chapter 2) when they revisited MMR designs twenty years after their important chapter was first published in 2003 tashakkori_handbook_2003 , the most recent studies decreased their focus on the quaN and quaL components and increased attention towards the integration processes across design types. To reflect this change about the central elements of the MMR designs, Figure 1 represents the simplification of the interconnection of the core components of a MMR design. From this perspective, treating the data beyond quaL and quaN analysis and gaining insight through integration and meta-inferences [^2] See Appendix for a glossary of terms used through the paper. Integration is discussed in detail later in the paper. become central.
@@
To conclude our brief overview of MMR, current trends in MMR can be listed as a call for openness and creativity during the research process, integrating mixed methods in interdisciplinary research, and the use of digital tools for data collection and analysis. Furthermore, researchers will be increasingly challenged to keep pace with a rapid evolution in design terminology and naming practices. Given such dynamism, paying attention to the many sources of contextual influences on the design of MMR becomes paramount poth_sage_2023 . Recognizing the challenges these trends impose on researchers, we provide practical design guidance of MMR for SE researchers in our paper. In the next section, we describe four fundamental principles to be followed when using MMR.
@@ 3 Principles to Guide Mixed Method Research
Building on the references discussed in Section 2, along with the MMR characteristics highlighted by Poth poth_innovation_2018 , Ladner ladner_mixed_2019 , Creswell and Plano Clark creswell_designing_2018 , we synthesize four guiding principles for conducting MMR in Software Engineering (see Figure 2). Our guiding principles are oriented by a pragmatic approach focused on describing key elements and properties of MMR rather than emphasizing the methodologies or philosophical foundations.
Two of the principles presented in our paper, methodological rationale and development of novel integrated insights (meta-inferences), relate to understanding why we should consider using a MMR design. The other two, procedural rigor and ethical research, relate to understanding how we should go about designing and applying MMR. While some aspects of these principles apply to single method applications and can be considered good research practice in general, others are particularly relevant to MMR and generally more complex in that context. We delve into these four principles below.
Figure 2: Four Guiding Principles for Mixed Method Research in Software Engineering.
@@ 3.1 Methodological Rationale
Principle of Methodological Rationale: Mixed method research study designs should explain how the mixed research design and data methods are relevant for addressing a given research problem. By adhering to this principle, an MMR study description can address the question why did we select a particular MMR research design and set of procedures?
There are several conceptual reasons for using mixed methods [^3] The reasons we list are inspired by or earlier proposed by Ladner ladner_mixed_2019 , Bryman bryman_research_2007 , Creswell and Plano Clark creswell_designing_2018 and McGrath mcgrath_methodology_1995 .. The rationale for the methodology may be one or more of the following:
  • Complementarity: one method helps to deepen, confirm or enhance the insights of another method ladner_mixed_2019 ; bryman_social_2012 .
  • Expansion: one method is used to ask different questions that are apparent up front or emerge throughout the study (in parallel or sequentially) ladner_mixed_2019 ; bryman_social_2012 .
  • Development: one method is used to inform or improve another method (e.g., interviews may be used to inform the development of a questionnaire) ladner_mixed_2019 ; bryman_social_2012 .
  • Triangulation: a different type of method is used to corroborate earlier insights leveraging multiple quality criteria afforded by use of different methods mcgrath_methodology_1995 notably generalizability, realism, precision, and control.
  • Credibility: another method may be used to increase the integrity and truth of the findings from a former method bryman_social_2012 . Mixed methods can increase credibility by bringing multi-dimensional philosophical and theoretical lenses to the findings mertens_mixed_2013 .
  • Explanation: one method can be used to resolve previous contradictory, surprising or inconclusive findings bryman_integrating_2006 .
@@ 3.2 Novel Integrated Insights
Principle of Novel Integrated Insights: Mixed methods research studies should generate new insights through the integration of the methods, which in turn addresses the reason(s) for the use of mixed methods. An MMR study report should clarify the novelty of the study findings and explain how methods and/or data were integrated. By adhering to this principle, an MMR study report can address the question what did we gain from using MMR?
An MMR study should deliver on its promise of additional benefits to what could have been afforded by using a single method poth_innovation_2018 . Typically, once data analysis for each quaL/quaN method is complete, the researcher compiles the results and findings, (a phase often referred to as the inferential stage teddlie_foundations_2009 ). In an MMR study, inferences are also drawn from both quaL and quaN analyses. However, to achieve a truly integrated approach, these conclusions should be synthesized into meta-inferences (see Fig. 1), which may involve generating additional insights or linking findings from both methods to broader theories or implications [^5] Venkatesh et al. venkatesh2013bridging provide an in-depth discussion on development and assessment of the quality of meta-inferences.. An MMR study may bring novel insights in terms of one or more of the following:
@@
  • Explaining unexpected results: where the use of one method produced unexpected results, another can be employed in a bid to further explore why those unexpected results were encountered.
@@ 3.3 Procedural Rigor
Principle of Procedural Rigor: Mixed method studies should apply the study methods rigorously and explain them effectively. By adhering to this principle, an MMR study report can address the question how well did we conduct the study?
Rigor in any research study is always a critical concern, and it is often a difficult concept for researchers to know how to assess. Rigor is particularly challenging for mixed methods. MMR studies should first adhere to the rigor of the individual studies, but there are also some concerns that apply to the “mixing” aspect. For this principle, we are guided by a mixed methods rigor appraisal tool called MMAT (Mixed Methods Appraisal Tool) oliveira_mixed_2021 . The following are posed as questions to be considered when self-reflecting on the rigor of a mixed design:
@@
  • Are the different methods used rigorously conducted? MMR studies should adhere to the rigor of the individual studies used in the mixed study (referring to the appropriate quality criteria for each study [^7] See the ACM SIGSOFT Empirical Standards for Software Engineering sigsoft . Other references could be consulted for further considerations on the aspects of quality and validation in the MMR integration process, such as Teddlie and Tashakkori teddlie_foundations_2009 and Venkatesh et al. venkatesh2013bridging ; venkatesh2016guidelines .).